### v2

@@ 3.4 The update of $\mathcal{L}$
Before we compute $\mathcal{L}^{k+1}$, we define the transpose of $\mathbf{R}$ by $\mathbf{R}^{\top}:$ $\mathbb{R}^{m_{1}\times m_{2}\times m_{3}\times N}\to\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ satisfying $\langle\mathbf{R}(\mathcal{L}),[\mathcal{Y}]\rangle=\langle\mathcal{L},\mathbf{R}^{\top}([\mathcal{Y}])\rangle$ for any $\mathcal{L}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, and $[\mathcal{Y}]\in\mathbb{R}^{m_{1}\times m_{2}\times m_{3}\times N}$. Then we have
$\|\mathbf{R}(\mathcal{L})\|_{F}^{2}=\langle\mathcal{L},\mathbf{R}^{\top}\mathbf{R}(\mathcal{L})\rangle_{F}=\langle\mathcal{L},\mathcal{W}_{\mathbf{R}}\odot\mathcal{L}\rangle_{F},$
where $\mathcal{W}_{\mathbf{R}}\in\mathbb{R}_{++}^{I_{1}\times I_{2}\times I_{3}}$ denotes a weight tensor such that the component-wise multiplication with $\mathcal{W}_{\mathbf{R}}$ is an equivalent operation of $\mathbf{R}^{\top}\mathbf{R}$, i.e., $\mathcal{W}_{\mathbf{R}}\odot\operatorname{Id}=\mathbf{R}^{\top}\mathbf{R}$, $\odot$ denotes the component-wise multiplication, and $\operatorname{Id}$ denotes the identity mapping on $\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$. Then $\mathcal{L}^{k+1}$ can be computed in a unique closed form as follows
(12) $\mathcal{L}^{k+1}=$ $(\delta\mathcal{W}_{\mathbf{R}}+\mathcal{I})^{-1}\odot\left(\delta\mathbf{R}^{\top}([\mathcal{Y}^{k+1}])+\mathcal{D}-\mathcal{S}^{k+1}\right),$
@@
(13) $[\mathcal{Y}^{k+1}]=[\mathcal{G}^{k+1}]\times_{1}[X_{1}^{k+1}]\times_{2}[X_{2}^{k+1}]\times_{3}[X_{3}^{k+1}],$
$\mathcal{I}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ denotes the tensor with all entries equal to $1$, and $\mathcal{W}_{\mathbf{R}}^{-1}$ denotes the component-wise inverse of $\mathcal{W}_{\mathbf{R}}$, i.e., the $(i_{1},i_{2},i_{3})$-th entry of $\mathcal{W}_{\mathbf{R}}^{-1}$ is equal to $1/(\mathcal{W}_{\mathbf{R}})_{i_{1}i_{2}i_{3}}$.
@@
1:  Initialize $(\mathcal{S}^{0},[{X}_{1}^{0}],[{X}_{2}^{0}],[{X}_{3}^{0}],[\mathcal{G}^{0}],\mathcal{L}^{0})$ with $[{X}_{i}^{0}]\in[\mathbb{S}_{m_{i},n_{i}}]$;
2:  Set the tensor extraction operator $\mathbf{R}$;
3:  Set parameters $\alpha_{\mathcal{S}},\alpha_{X},\alpha_{\mathcal{G}}>0$;
@@
11:  $(\mathcal{S}^{k},[{X}_{1}^{k}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})$.
@@ 4 Convergence Analysis of The P-BCD Algorithm
Let $\mathcal{Z}:=(\mathcal{S},[X_{1}],[X_{2}],[X_{3}],[\mathcal{G}],\mathcal{L})$. First, we define the first-order optimality condition of the orthogonal constrained optimization problem as in (1). The point $\bar{\mathcal{Z}}:=(\bar{\mathcal{S}},[\bar{X}_{1}],[\bar{X}_{2}],[\bar{X}_{3}],[\bar{\mathcal{G}}],\bar{\mathcal{L}})$ is a first-order stationary point of problem (1) if $0\in\partial\Phi(\bar{\mathcal{Z}})$, that is,
${0}\in\bar{\mathcal{L}}+\bar{\mathcal{S}}-\mathcal{D}+\gamma\partial\varphi(\bar{\mathcal{S}}),$
(14) ${0}=\operatorname{grad}_{[{X}_{i}]}H([\bar{X}_{1}],[\bar{X}_{2}],[\bar{X}_{3}],[\bar{\mathcal{G}}],\bar{\mathcal{L}}),\quad[\bar{X}_{i}]^{\top}[\bar{X}_{i}]=[\mathbf{I}_{n_{i}}],i=1,2,3,$
${0}\in\delta\nabla_{[\mathcal{G}]}H([\bar{X}_{1}],[\bar{X}_{2}],[\bar{X}_{3}],[\bar{\mathcal{G}}],\bar{\mathcal{L}})+\partial\psi([\bar{\mathcal{G}}]),$
${0}=\bar{\mathcal{L}}+\bar{\mathcal{S}}-\mathcal{D}+\delta\nabla_{\mathcal{L}}H([\bar{X}_{1}],[\bar{X}_{2}],[\bar{X}_{3}],[\bar{\mathcal{G}}],\bar{\mathcal{L}}),$
where $\operatorname{grad}_{[{X}_{i}]}H([\bar{X}_{1}],[\bar{X}_{2}],[\bar{X}_{3}],[\bar{\mathcal{G}}],\bar{\mathcal{L}})$ denotes the Riemannian gradient of $H$ with respect to $[{X}_{i}]$ evaluated at $([\bar{X}_{1}],[\bar{X}_{2}],[\bar{X}_{3}],[\bar{\mathcal{G}}],\bar{\mathcal{L}})$, $i=1,2,3$, and $\partial\varphi$ and $\partial\psi$ denote the subdifferentials of $\varphi$ and $\psi$, respectively. Then we compute the gradients of $H$ explicitly and replace the optimality condition for orthogonal constraints as in (14) using an equivalent condition introduced in [17]. Hence, we call $\bar{\mathcal{Z}}$ is a first-order stationary point of problem (1) if
(15a) ${0}\in\bar{\mathcal{L}}+\bar{\mathcal{S}}-\mathcal{D}+\gamma\partial\varphi(\bar{\mathcal{S}}),$
(15b) ${0}=([\mathbf{I}_{m_{i}}]-[\bar{X}_{i}][\bar{X}_{i}]^{\top})[\bar{H}_{i}],$
(15c) ${0}=[\bar{H}_{i}]^{\top}[\bar{X}_{i}]-[\bar{X}_{i}]^{\top}[\bar{H}_{i}],$
(15d) $[\bar{X}_{i}]^{\top}[\bar{X}_{i}]=[\mathbf{I}_{n_{i}}],$
(15e) ${0}\in\delta([\bar{\mathcal{G}}]-[\bar{\mathcal{O}}])+\partial\psi([\bar{\mathcal{G}}]),$
(15f) ${0}=\bar{\mathcal{L}}+\bar{\mathcal{S}}-\mathcal{D}+\delta\mathcal{W}_{\mathbf{R}}\odot\bar{\mathcal{L}}-\delta\mathbf{R}^{\top}([\bar{\mathcal{Y}}]),$
@@
$[\bar{P}_{i}]=$ $(\mathbf{R}(\bar{\mathcal{L}}))_{[(i)]}$
$[\bar{Q}_{1}]=$ $([\bar{\mathcal{G}}]\times_{2}[\bar{X}_{2}]\times_{3}[\bar{X}_{3}])_{[(1)]}$
$[\bar{Q}_{2}]=$ $([\bar{\mathcal{G}}]\times_{1}[\bar{X}_{1}]\times_{3}[\bar{X}_{3}])_{[(2)]}$
$[\bar{Q}_{3}]=$ $([\bar{\mathcal{G}}]\times_{1}[\bar{X}_{1}]\times_{2}[\bar{X}_{2}])_{[(3)]}$
$[\bar{H}_{i}]=$ $([\bar{X}_{i}][\bar{Q}_{i}]-[\bar{P}_{i}])[\bar{Q}_{i}]^{\top}$
$[\bar{\mathcal{O}}]=$ $\mathbf{R}(\bar{\mathcal{L}})\times_{1}[\bar{X}_{1}]^{\top}\times_{2}[\bar{X}_{2}]^{\top}\times_{3}[\bar{X}_{3}]^{\top}$
$[\bar{\mathcal{Y}}]=$ $[\bar{\mathcal{G}}]\times_{1}[\bar{X}_{1}]\times_{2}[\bar{X}_{2}]\times_{3}[\bar{X}_{3}].$
@@
    (16) $\begin{split}&\Phi(\mathcal{Z}^{k})-\Phi(\mathcal{Z}^{k+1})\\ \geq&\frac{\alpha}{2}\|\mathcal{S}^{k+1}-\mathcal{S}^{k}\|^{2}_{F}+\frac{\alpha}{2}\sum_{i=1}^{3}\|[{X}_{i}^{k+1}]-[{X}_{i}^{k}]\|^{2}_{F}+\frac{\alpha}{2}\|[\mathcal{G}^{k+1}]-[\mathcal{G}^{k}]\|^{2}_{F}.\end{split}$
@@
    $\lim_{k\to\infty}\|\mathcal{S}^{k+1}-\mathcal{S}^{k}\|_{F}=0$, $\lim_{k\to\infty}\|[{X}_{i}^{k+1}]-[{X}_{i}^{k}]\|_{F}=0$, for $i=1,2,3$, and $\lim_{k\to\infty}\|[\mathcal{G}^{k+1}]-[\mathcal{G}^{k}]\|_{F}=0$.
@@
$\Phi(\mathcal{Z}^{k})-\Phi(\mathcal{S}^{k+1},[{X}_{1}^{k}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})$
@@
where $\alpha>0$. Next, it follows from the update of $[X_{i}]$ and Lemma 3.1 that
$\Phi(\mathcal{S}^{k+1},[{X}_{1}^{k}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-\Phi(\mathcal{S}^{k+1},[{X}_{1}^{k+1}],[{X}_{2}^{k+1}],[{X}_{3}^{k+1}],[\mathcal{G}^{k}],\mathcal{L}^{k})$
$\geq$ $\frac{\alpha_{X}}{2}\sum_{i=1}^{3}\|[{X}_{i}^{k+1}]-[{X}_{i}^{k}]\|^{2}_{F}.$
@@
$\Phi(\mathcal{S}^{k+1},[{X}_{1}^{k+1}],[{X}_{2}^{k+1}],[{X}_{3}^{k+1}],[\mathcal{G}^{k}],\mathcal{L}^{k})$
$~{}~{}~{}~{}~{}~{}~{}-\Phi(\mathcal{S}^{k+1},[{X}_{1}^{k+1}],[{X}_{2}^{k+1}],[{X}_{3}^{k+1}],[\mathcal{G}^{k+1}],\mathcal{L}^{k})\geq\frac{\alpha_{\mathcal{G}}}{2}\|[\mathcal{G}^{k+1}]-[\mathcal{G}^{k}]\|^{2}_{F}$
@@
$\Phi(\mathcal{S}^{k+1},[{X}_{1}^{k+1}],[{X}_{2}^{k+1}],[{X}_{3}^{k+1}],[\mathcal{G}^{k+1}],\mathcal{L}^{k})-\Phi(\mathcal{Z}^{k+1})\geq 0.$
Combining the inequalities above, we obtain (16) with $\alpha=\min\{\alpha_{\mathcal{S}},\alpha_{\mathcal{G}},\alpha_{X}\}>0$.
(ii) Since $[X_{i}^{k}]^{\top}[X_{i}^{k}]=[\mathbf{I}_{n_{i}}]$ for each $i=1,2,3$, we have the sequence $\{[X_{i}^{k}]\}$ is bounded. By (i), we have $\Phi(\mathcal{Z}^{k})\leq\Phi(\mathcal{Z}^{0})$. Also, we observe that $\Phi(\mathcal{Z}^{k})\geq\gamma\varphi(\mathcal{S}^{k})+\psi([\mathcal{G}^{k}])\geq 0$. Since
@@
we must have the sequences $\{\mathcal{S}^{k}\}$ and $\{[\mathcal{G}^{k}]\}$ are bounded. As shown in (12) that $\mathcal{L}^{k}$ is uniquely determined by $\mathcal{S}^{k},[{X}_{1}^{k}],[{X}_{2}^{k}],[{X}_{3}^{k}]$ and $[\mathcal{G}^{k}]$, the sequence $\{\mathcal{L}^{k}\}$ is also bounded.
@@
$\sum^{K-1}_{k=0}\|\mathcal{S}^{k+1}-\mathcal{S}^{k}\|^{2}_{F}+\sum^{K-1}_{k=0}\sum_{i=1}^{3}\|[{X}_{i}^{k+1}]-[{X}_{i}^{k}]\|^{2}_{F}+\sum^{K-1}_{k=0}\|[\mathcal{G}^{k+1}]-[\mathcal{G}^{k}]\|^{2}_{F}$
@@
Taking the limits of both sides of the inequality as $K\rightarrow\infty$, we have $\sum^{\infty}_{k=0}\|\mathcal{S}^{k+1}-\mathcal{S}^{k}\|^{2}_{F}<\infty$, $\sum_{k=0}^{\infty}\|[{X}_{i}^{k+1}]-[{X}_{i}^{k}]\|^{2}_{F}<\infty$ and $\sum_{k=0}^{\infty}\|[\mathcal{G}^{k+1}]-[\mathcal{G}^{k}]\|^{2}_{F}<\infty$. Then assertion (iii) immediately holds.
@@
Since $\|[\mathcal{G}^{k+1}]\times_{i}[X_{i}]-[\mathcal{G}^{k}]\times_{i}[X_{i}]\|_{F}=\|[\mathcal{G}^{k+1}]-[\mathcal{G}^{k}]\|_{F}$ for any $[X_{i}]\in[\mathbb{S}_{m_{i},n_{i}}]$ and $\|[\mathcal{G}]\times_{i}[X_{i}^{k+1}]-[\mathcal{G}]\times_{i}[X_{i}^{k}]\|_{F}\leq\|[\mathcal{G}]\|_{F}\|[X_{i}^{k+1}]-[X_{i}^{k}]\|_{F}$, we have
$\|[\mathcal{Y}^{k+1}]-[\mathcal{Y}^{k}]\|_{F}\leq$ $\|[\mathcal{G}^{k+1}]-[\mathcal{G}^{k}]\|_{F}+c_{1}\|[{X}_{1}^{k+1}]-[{X}_{1}^{k}]\|_{F}$
$+c_{1}\|[{X}_{2}^{k+1}]-[{X}_{2}^{k}]\|_{F}+c_{1}\|[{X}_{3}^{k+1}]-[{X}_{3}^{k}]\|_{F},$
@@
where $c_{2}=\max(\mathcal{W}_{\mathbf{R}})_{i_{1}i_{2}i_{3}}$. Then by assertion (iii) in Theorem 4.1, we can have $\lim_{k\to\infty}\|\mathcal{L}^{k+1}-\mathcal{L}^{k}\|_{F}=0$.
Third, we apply the results in [17] to the updates of $[X_{i}]$ in our proposed algorithm. Those results are useful for proving three equalities of substationarity, symmetry, and feasibility for $[X_{i}]$. According to Lemma 3.3 in [17], we can have the following lemma and then we prove the decrease of the function value $H$ after each update of $[X_{i}]$.

### v3

@@ 3.4 The update of $\mathcal{L}$
Before we compute $\mathcal{L}^{k+1}$, we define the transpose of $\mathscr{R}$ by $\mathscr{R}^{\top}:$ $\mathbb{R}^{m_{1}\times m_{2}\times m_{3}\times N}\to\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ satisfying $\langle\mathscr{R}(\mathcal{L}),[\mathcal{Y}]\rangle=\langle\mathcal{L},\mathscr{R}^{\top}([\mathcal{Y}])\rangle$ for any $\mathcal{L}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, and $[\mathcal{Y}]\in\mathbb{R}^{m_{1}\times m_{2}\times m_{3}\times N}$. Then we have
$\|\mathscr{R}(\mathcal{L})\|_{F}^{2}=\langle\mathcal{L},\mathscr{R}^{\top}\mathscr{R}(\mathcal{L})\rangle_{F}=\langle\mathcal{L},\mathcal{W}_{\mathscr{R}}\odot\mathcal{L}\rangle_{F},$
where $\mathcal{W}_{\mathscr{R}}\in\mathbb{R}_{++}^{I_{1}\times I_{2}\times I_{3}}$ denotes a weight tensor such that the component-wise multiplication with $\mathcal{W}_{\mathscr{R}}$ is an equivalent operation of $\mathscr{R}^{\top}\mathscr{R}$, i.e., $\mathcal{W}_{\mathscr{R}}\odot\operatorname{Id}=\mathscr{R}^{\top}\mathscr{R}$, $\odot$ denotes the component-wise multiplication, and $\operatorname{Id}$ denotes the identity mapping on $\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$. Then $\mathcal{L}^{k+1}$ can be computed in a unique closed form as follows
(13) $\mathcal{L}^{k+1}=$ $(\delta\mathcal{W}_{\mathscr{R}}+\mathcal{I})^{-1}\odot\left(\delta\mathscr{R}^{\top}([\mathcal{Y}^{k+1}])+\mathcal{D}-\mathcal{S}^{k+1}\right),$
@@
(14) $[\mathcal{Y}^{k+1}]=[\mathcal{G}^{k+1}]\times_{1}[\boldsymbol{X}_{1}^{k+1}]\times_{2}[\boldsymbol{X}_{2}^{k+1}]\times_{3}[\boldsymbol{X}_{3}^{k+1}],$
$\mathcal{I}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ denotes the tensor with all entries equal to $1$, and $\mathcal{W}_{\mathscr{R}}^{-1}$ denotes the component-wise inverse of $\mathcal{W}_{\mathscr{R}}$, i.e., the $(i_{1},i_{2},i_{3})$-th entry of $\mathcal{W}_{\mathscr{R}}^{-1}$ is equal to $1/(\mathcal{W}_{\mathscr{R}})_{i_{1}i_{2}i_{3}}$.
@@
1:  Initialize $(\mathcal{S}^{0},[\boldsymbol{X}_{1}^{0}],[\boldsymbol{X}_{2}^{0}],[\boldsymbol{X}_{3}^{0}],[\mathcal{G}^{0}],\mathcal{L}^{0})$ with $[\boldsymbol{X}_{i}^{0}]\in[\mathbb{S}_{m_{i},n_{i}}]$;
2:  Set the tensor extraction operator $\mathscr{R}$;
3:  Set parameters $\alpha_{\mathcal{S}},\alpha_{\boldsymbol{X}},\alpha_{\mathcal{G}}>0$;
@@
11:  $(\mathcal{S}^{k},[\boldsymbol{X}_{1}^{k}],[\boldsymbol{X}_{2}^{k}],[\boldsymbol{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})$.
@@ 4 Convergence Analysis of The P-BCD Algorithm
Let $\mathcal{Z}:=(\mathcal{S},[\boldsymbol{X}_{1}],[\boldsymbol{X}_{2}],[\boldsymbol{X}_{3}],[\mathcal{G}],\mathcal{L})$. First, we define the first-order optimality condition of the orthogonal constrained optimization problem as in (1). The point $\bar{\mathcal{Z}}:=(\bar{\mathcal{S}},[\bar{\boldsymbol{X}}_{1}],[\bar{\boldsymbol{X}}_{2}],[\bar{\boldsymbol{X}}_{3}],[\bar{\mathcal{G}}],\bar{\mathcal{L}})$ is a first-order stationary point of problem (1) if $\boldsymbol{0}\in\partial\Phi(\bar{\mathcal{Z}})$, that is,
$\boldsymbol{0}\in\bar{\mathcal{L}}+\bar{\mathcal{S}}-\mathcal{D}+\gamma\partial\varphi(\bar{\mathcal{S}}),$
(15) $\boldsymbol{0}=\operatorname{grad}_{[\boldsymbol{X}_{i}]}H([\bar{\boldsymbol{X}}_{1}],[\bar{\boldsymbol{X}}_{2}],[\bar{\boldsymbol{X}}_{3}],[\bar{\mathcal{G}}],\bar{\mathcal{L}}),\quad[\bar{\boldsymbol{X}}_{i}]^{\top}[\bar{\boldsymbol{X}}_{i}]=[\boldsymbol{I}_{n_{i}}],i=1,2,3,$
$\boldsymbol{0}\in\delta\nabla_{[\mathcal{G}]}H([\bar{\boldsymbol{X}}_{1}],[\bar{\boldsymbol{X}}_{2}],[\bar{\boldsymbol{X}}_{3}],[\bar{\mathcal{G}}],\bar{\mathcal{L}})+\partial\psi([\bar{\mathcal{G}}]),$
$\boldsymbol{0}=\bar{\mathcal{L}}+\bar{\mathcal{S}}-\mathcal{D}+\delta\nabla_{\mathcal{L}}H([\bar{\boldsymbol{X}}_{1}],[\bar{\boldsymbol{X}}_{2}],[\bar{\boldsymbol{X}}_{3}],[\bar{\mathcal{G}}],\bar{\mathcal{L}}),$
where $\operatorname{grad}_{[\boldsymbol{X}_{i}]}H([\bar{\boldsymbol{X}}_{1}],[\bar{\boldsymbol{X}}_{2}],[\bar{\boldsymbol{X}}_{3}],[\bar{\mathcal{G}}],\bar{\mathcal{L}})$ denotes the Riemannian gradient of $H$ with respect to $[\boldsymbol{X}_{i}]$ evaluated at $([\bar{\boldsymbol{X}}_{1}],[\bar{\boldsymbol{X}}_{2}],[\bar{\boldsymbol{X}}_{3}],[\bar{\mathcal{G}}],\bar{\mathcal{L}})$, $i=1,2,3$, and $\partial\varphi$ and $\partial\psi$ denote the subdifferentials of $\varphi$ and $\psi$, respectively. Then we compute the gradients of $H$ explicitly and replace the optimality condition for orthogonal constraints as in (15) using an equivalent condition introduced in [17]. Hence, we call $\bar{\mathcal{Z}}$ is a first-order stationary point of problem (1) if
(16a) $\boldsymbol{0}\in\bar{\mathcal{L}}+\bar{\mathcal{S}}-\mathcal{D}+\gamma\partial\varphi(\bar{\mathcal{S}}),$
(16b) $\boldsymbol{0}=([\boldsymbol{I}_{m_{i}}]-[\bar{\boldsymbol{X}}_{i}][\bar{\boldsymbol{X}}_{i}]^{\top})[\bar{\boldsymbol{H}}_{i}],$
(16c) $\boldsymbol{0}=[\bar{\boldsymbol{H}}_{i}]^{\top}[\bar{\boldsymbol{X}}_{i}]-[\bar{\boldsymbol{X}}_{i}]^{\top}[\bar{\boldsymbol{H}}_{i}],$
(16d) $[\bar{\boldsymbol{X}}_{i}]^{\top}[\bar{\boldsymbol{X}}_{i}]=[\mathbf{I}_{n_{i}}],$
(16e) $\boldsymbol{0}\in\delta([\bar{\mathcal{G}}]-[\bar{\mathcal{O}}])+\partial\psi([\bar{\mathcal{G}}]),$
(16f) $\boldsymbol{0}=\bar{\mathcal{L}}+\bar{\mathcal{S}}-\mathcal{D}+\delta\mathcal{W}_{\mathscr{R}}\odot\bar{\mathcal{L}}-\delta\mathscr{R}^{\top}([\bar{\mathcal{Y}}]),$
@@
$[\bar{\boldsymbol{P}}_{i}]=$ ((¯$(\mathscr{R}(\bar{\mathcal{L}}))_{[(i)]}$
$[\bar{\boldsymbol{Q}}_{1}]=$ $([\bar{\mathcal{G}}]\times_{2}[\bar{\boldsymbol{X}}_{2}]\times_{3}[\bar{\boldsymbol{X}}_{3}])_{[(1)]}$
$[\bar{\boldsymbol{Q}}_{2}]=$ $([\bar{\mathcal{G}}]\times_{1}[\bar{\boldsymbol{X}}_{1}]\times_{3}[\bar{\boldsymbol{X}}_{3}])_{[(2)]}$
$[\bar{\boldsymbol{Q}}_{3}]=$ $([\bar{\mathcal{G}}]\times_{1}[\bar{\boldsymbol{X}}_{1}]\times_{2}[\bar{\boldsymbol{X}}_{2}])_{[(3)]}$
$[\bar{\boldsymbol{H}}_{i}]=$ $([\bar{\boldsymbol{X}}_{i}][\bar{\boldsymbol{Q}}_{i}]-[\bar{\boldsymbol{P}}_{i}])[\bar{\boldsymbol{Q}}_{i}]^{\top}$
$[\bar{\mathcal{O}}]=$ (¯)×$\mathscr{R}(\bar{\mathcal{L}})\times_{1}[\bar{\boldsymbol{X}}_{1}]^{\top}\times_{2}[\bar{\boldsymbol{X}}_{2}]^{\top}\times_{3}[\bar{\boldsymbol{X}}_{3}]^{\top}$
$[\bar{\mathcal{Y}}]=$ $[\bar{\mathcal{G}}]\times_{1}[\bar{\boldsymbol{X}}_{1}]\times_{2}[\bar{\boldsymbol{X}}_{2}]\times_{3}[\bar{\boldsymbol{X}}_{3}].$
@@
    (17) $\begin{split}&\Phi(\mathcal{Z}^{k})-\Phi(\mathcal{Z}^{k+1})\\ \geq&\frac{\alpha}{2}\|\mathcal{S}^{k+1}-\mathcal{S}^{k}\|^{2}_{F}+\frac{\alpha}{2}\sum_{i=1}^{3}\|[\boldsymbol{X}_{i}^{k+1}]-[\boldsymbol{X}_{i}^{k}]\|^{2}_{F}+\frac{\alpha}{2}\|[\mathcal{G}^{k+1}]-[\mathcal{G}^{k}]\|^{2}_{F}.\end{split}$
@@
    $\lim_{k\to\infty}\|\mathcal{S}^{k+1}-\mathcal{S}^{k}\|_{F}=0$, $\lim_{k\to\infty}\|[\boldsymbol{X}_{i}^{k+1}]-[\boldsymbol{X}_{i}^{k}]\|_{F}=0$, for $i=1,2,3$, and $\lim_{k\to\infty}\|[\mathcal{G}^{k+1}]-[\mathcal{G}^{k}]\|_{F}=0$.
@@
$\Phi(\mathcal{Z}^{k})-\Phi(\mathcal{S}^{k+1},[\boldsymbol{X}_{1}^{k}],[\boldsymbol{X}_{2}^{k}],[\boldsymbol{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})$
@@
where $\alpha>0$. Next, it follows from the update of $[\boldsymbol{X}_{i}]$ and Lemma 3.1 that
$\Phi(\mathcal{S}^{k+1},[\boldsymbol{X}_{1}^{k}],[\boldsymbol{X}_{2}^{k}],[\boldsymbol{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-\Phi(\mathcal{S}^{k+1},[\boldsymbol{X}_{1}^{k+1}],[\boldsymbol{X}_{2}^{k+1}],[\boldsymbol{X}_{3}^{k+1}],[\mathcal{G}^{k}],\mathcal{L}^{k})$
$\geq$ $\frac{\alpha_{\boldsymbol{X}}}{2}\sum_{i=1}^{3}\|[\boldsymbol{X}_{i}^{k+1}]-[\boldsymbol{X}_{i}^{k}]\|^{2}_{F}.$
@@
$\Phi(\mathcal{S}^{k+1},[\boldsymbol{X}_{1}^{k+1}],[\boldsymbol{X}_{2}^{k+1}],[\boldsymbol{X}_{3}^{k+1}],[\mathcal{G}^{k}],\mathcal{L}^{k})$
$~{}~{}~{}~{}~{}~{}~{}-\Phi(\mathcal{S}^{k+1},[\boldsymbol{X}_{1}^{k+1}],[\boldsymbol{X}_{2}^{k+1}],[\boldsymbol{X}_{3}^{k+1}],[\mathcal{G}^{k+1}],\mathcal{L}^{k})\geq\frac{\alpha_{\mathcal{G}}}{2}\|[\mathcal{G}^{k+1}]-[\mathcal{G}^{k}]\|^{2}_{F}$
@@
$\Phi(\mathcal{S}^{k+1},[\boldsymbol{X}_{1}^{k+1}],[\boldsymbol{X}_{2}^{k+1}],[\boldsymbol{X}_{3}^{k+1}],[\mathcal{G}^{k+1}],\mathcal{L}^{k})-\Phi(\mathcal{Z}^{k+1})\geq 0.$
Combining the inequalities above, we obtain (17) with $\alpha=\min\{\alpha_{\mathcal{S}},\alpha_{\mathcal{G}},\alpha_{\boldsymbol{X}}\}>0$.
(ii) Since $[\boldsymbol{X}_{i}^{k}]^{\top}[\boldsymbol{X}_{i}^{k}]=[\boldsymbol{I}_{n_{i}}]$ for each $i=1,2,3$, we have the sequence $\{[\boldsymbol{X}_{i}^{k}]\}$ is bounded. By (i), we have $\Phi(\mathcal{Z}^{k})\leq\Phi(\mathcal{Z}^{0})$. Also, we observe that $\Phi(\mathcal{Z}^{k})\geq\gamma\varphi(\mathcal{S}^{k})+\psi([\mathcal{G}^{k}])\geq 0$. Since
@@
we must have the sequences $\{\mathcal{S}^{k}\}$ and $\{[\mathcal{G}^{k}]\}$ are bounded. As shown in (13) that $\mathcal{L}^{k}$ is uniquely determined by $\mathcal{S}^{k},[\boldsymbol{X}_{1}^{k}],[\boldsymbol{X}_{2}^{k}],[\boldsymbol{X}_{3}^{k}]$ and $[\mathcal{G}^{k}]$, the sequence $\{\mathcal{L}^{k}\}$ is also bounded.
@@
$\sum^{K-1}_{k=0}\|\mathcal{S}^{k+1}-\mathcal{S}^{k}\|^{2}_{F}+\sum^{K-1}_{k=0}\sum_{i=1}^{3}\|[\boldsymbol{X}_{i}^{k+1}]-[\boldsymbol{X}_{i}^{k}]\|^{2}_{F}+\sum^{K-1}_{k=0}\|[\mathcal{G}^{k+1}]-[\mathcal{G}^{k}]\|^{2}_{F}$
@@
Taking the limits of both sides of the inequality as $K\rightarrow\infty$, we have $\sum^{\infty}_{k=0}\|\mathcal{S}^{k+1}-\mathcal{S}^{k}\|^{2}_{F}<\infty$, $\sum_{k=0}^{\infty}\|[\boldsymbol{X}_{i}^{k+1}]-[\boldsymbol{X}_{i}^{k}]\|^{2}_{F}<\infty$ and $\sum_{k=0}^{\infty}\|[\mathcal{G}^{k+1}]-[\mathcal{G}^{k}]\|^{2}_{F}<\infty$. Then assertion (iii) immediately holds.
@@
Since $\|[\mathcal{G}^{k+1}]\times_{i}[\boldsymbol{X}_{i}]-[\mathcal{G}^{k}]\times_{i}[\boldsymbol{X}_{i}]\|_{F}=\|[\mathcal{G}^{k+1}]-[\mathcal{G}^{k}]\|_{F}$ for any $[\boldsymbol{X}_{i}]\in[\mathbb{S}_{m_{i},n_{i}}]$ and $\|[\mathcal{G}]\times_{i}[\boldsymbol{X}_{i}^{k+1}]-[\mathcal{G}]\times_{i}[\boldsymbol{X}_{i}^{k}]\|_{F}\leq\|[\mathcal{G}]\|_{F}\|[\boldsymbol{X}_{i}^{k+1}]-[\boldsymbol{X}_{i}^{k}]\|_{F}$, we have
$\|[\mathcal{Y}^{k+1}]-[\mathcal{Y}^{k}]\|_{F}\leq$ $\|[\mathcal{G}^{k+1}]-[\mathcal{G}^{k}]\|_{F}+c_{1}\|[\boldsymbol{X}_{1}^{k+1}]-[\boldsymbol{X}_{1}^{k}]\|_{F}$
$+c_{1}\|[\boldsymbol{X}_{2}^{k+1}]-[\boldsymbol{X}_{2}^{k}]\|_{F}+c_{1}\|[\boldsymbol{X}_{3}^{k+1}]-[\boldsymbol{X}_{3}^{k}]\|_{F},$
@@
where $c_{2}=\max(\mathcal{W}_{\mathscr{R}})_{i_{1}i_{2}i_{3}}$. Then by assertion (iii) in Theorem 4.1, we can have $\lim_{k\to\infty}\|\mathcal{L}^{k+1}-\mathcal{L}^{k}\|_{F}=0$.
Third, we apply the results in [17] to the updates of $[\boldsymbol{X}_{i}]$ in our proposed algorithm. Those results are useful for proving three equalities of substationarity, symmetry, and feasibility for $[\boldsymbol{X}_{i}]$. According to Lemma 3.3 in [17], we can have the following lemma and then we prove the decrease of the function value $H$ after each update of $[\boldsymbol{X}_{i}]$.